Determine a such that x2-11x+a and x2-14x +2a may have a common factor.

Grade:11

1 Answers

Vasantha Kumari
9 years ago

(x–b)(x–c) = x2–11x+a …(1)

b + c = 11

bc=a

(x–a)(x–d)=x2–14x+2a …(2)

b + d = 14

bd = 2a

Solving (1) and (2) equations,

B=8, c =3, d=6 and a = 24.

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty